# Bringing Generative AI to Adaptive Learning in Education

### v2

@@ 4 Bringing GenAI to Adaptive Learning
The recent achievements made by GenAI in the areas such as medicine (Zhang & Kamel Boulos, 2023), programming (Sarkar et al., 2022), and designing (Cai et al., 2023) underscore its exceptional potential in applications compared to the prior machine learning methods. In the following, we will outline the advantages GenAI that can offer to adaptive learning, along with examples of industrial applications. Then, we will provide insights from an educational perspective, delving into the disadvantages and challenges faced by the new framework.
@@ 4.1.1 Diversity and Dynamics
Integrating generative AI into adaptive learning offers significant benefits, notably its ability to produce dynamic and diverse outputs. Current ML-based adaptive learning systems, despite their automation (Wang et al., 2023b), still fall short of replicating the nuanced interactions of a skilled human teacher. GenAI can bridge this gap by generating real-time, tailored responses and learning plans based on each student’s immediate questions and learning status. Moreover, the recent advancements in large language models for context understanding (Brown et al., 2020) further enhance the potential for providing personalized feedback, offering a substantial leap in optimizing learning outcomes. Apart from providing new features, GenAI holds the potential to enhance current ML models in AL systems. As detailed in Section 3.1, knowledge tracing methods typically use learner’s question practice records, but these are often pre-set and designed for mainstream students, potentially hindering accurate assessment of individual learning status. GenAI, with its dynamic and diverse capabilities, can generate questions tailored to each learner’s past performance, aiding in more effective knowledge tracing. Furthermore, GenAI’s ability to diversify content, such as in math word problems (MWP) (Zhou et al., 2023b), reduces repetition and improves the accuracy of knowledge proficiency assessments. This adaptability makes GenAI a valuable asset in evolving and refining adaptive learning systems.
@@ 4.1.2 Multi-modality Convenience
As the strong capabilities of GenAI in dealing with multi-modal data have been verified in multiple studies (Bubeck et al., 2023), applying GenAI will enable adaptive learning to understand student’s learning status from a broader perspective. For example, the current accepted data format to adaptive learning methods are mainly limited to structured text, by integrating GenAI, the AL systems will be able to interpret leaner emotions in speech and facial expressions together with the text input. The enhanced input information will help the following analysis of the student’s learning characteristics such as learning style and learning status (Giannakos & Cukurova, 2023), which will also benefit the precise adaptions to the student’s learning paths. Meanwhile, as GenAI’s outputs can also be multi-modal (Wu et al., 2023), the generated learning materials for knowledge concepts can be more inspiring. For example, with the help of image generation models, the interpretations of geometry problems could be presented with figures or even a short animation, which will be easier for learners to understand compared to prior pure-text responses.
4.1.3 Broad General Knowledges
One of the common challenges for machine learning methods in education is the scarcity of high-quality annotated data, which is caused by its strong privacy properties and high annotation costs. Due to this fact, many of prior models suffer from over-fitting and poor generalization problems. Fortunately, by introducing GenAI, the above issues can be significantly alleviated as GenAI models, especially LLMs, are general-purpose foundation models (Brühl, 2023). Contributing to the broad general knowledge learned from the enormous sized pre-training data, GenAI has been widely recognized as a suitable solution in dealing with data limitation problems (Brown et al., 2020). Additionally, with the recently proposed few-shot learning methods (Dang et al., 2022), the adaptive learning with GenAI is expected to achieve great improvements in performance compared with the prior ones, which are trained from scratches. At last, examples of GenAI models, such as AlphaZero (Silver et al., 2017), and AlphaFold (Senior et al., 2020), in generating unconventional solutions suggest that GenAI may understand and utilize its learned knowledge in ways that are different from our human beings. By bringing GenAI to adaptive learning, we expect to see novel results, e.g., unexplored learning paths, during the usages. With such ”jump out of box” thoughts, we may get chances to explore more improvements to adaptive learning in the future.
4.2 Industrial Practice
Recognizing the benefits, leading institutions in adaptive learning research have begun to experiment with integrating large language models into their existing educational frameworks. In Table 1 of the Appendix, we list some of these pioneering institutions as representative examples of this trend. The majority of these business cases focus on leveraging GenAI to empower their original products and creating personalized learning experience to their customers.
@@ 4.3.1 Hallucination
The biggest issue with generative AI currently is its potential to produce content that doesn’t actually exist, a phenomenon known as ”Hallucination” (Ji et al., 2023). Hallucinations happened even in the most advanced generative AI models. Research suggests that educating students about the limitations of using Generative AI is crucial, emphasizing its susceptibility to errors when working with it (Strzelecki & ElArabawy, 2024; Bearman & Ajjawi, 2023). This brings forth numerous opportunities and research questions for the future, such as how to design the adaptive learning system to help students cultivate an appropriate reliance on AI-generated content (Vereschak et al., 2021). Furthermore, hallucination could become a new adaptive teaching method, such as creating new and possibly erroneous educational content and asking students to practice the role of a tutor in correcting AI’s mistakes (Ma et al., 2023).
@@ 4.3.2 Capability Decay
The ease and efficiency of obtaining information from GenAI can create a dependency, where students might prefer quick GenAI answers over the process of learning and discovery. This overreliance could erode critical thinking skills and reduce the students’ capacity for self-driven inquiry, as they might become accustomed to receiving immediate answers without engaging in the cognitive effort required for learning. One potential solution is to integrate generative AI as a tool that prompts further questioning and exploration rather than providing definitive answers. Encouraging projects and assignments that require independent research and critical analysis can also help maintain a balance between utilizing generative AI and fostering self-reliance. However, these suggestions only begin to address the broader challenge of integrating generative AI into education without compromising the development of independent learning skills. This issue remains an open question, presenting a wide array of opportunities for educators, technologists, and policymakers to explore innovative solutions that balance the benefits of generative AI with the imperative of nurturing autonomous, curious learners.
@@ 4.3.3 Fairness
Fairness is a crucial ethical concern, particularly in education where generative AI technologies like GPT models are being integrated (Fenu et al., 2022). These advancements, while beneficial, bring to light issues of equitable access and unbiased treatment among learners. A notable concern is the varied exposure of learners to different levels of generative AI. For instance, discrepancies between those accessing advanced systems like GPT-4 and those with older versions like GPT-3.5 can lead to unequal knowledge distribution, potentially exacerbating educational gaps. The effective deployment of generative AI in education also hinges on user familiarity and accessibility. Students well-acquainted with AI technologies stand to gain more, while those less experienced may encounter challenges and a steeper learning curve. Introducing standardized training programs might help bridge this gap, though such measures require meticulous implementation and oversight. Additionally, inherent biases within GenAI pose a significant challenge. These systems, limited by the biases in their training data, might offer skewed guidance, making vigilant monitoring essential to ensure fairness. In adaptive learning, where generative AI tailors instructional guidance, the risk of bias amplifies. Personalizing learning experiences is beneficial but needs to be weighed against the risk of unfairness. Ensuring an optimal, equitable, and unbiased educational experience for every learner involves navigating these complexities, making the balance between personalized learning and fairness both a technical hurdle and an ethical necessity in the integration of generative AI into educational systems.
@@ 4.3.4 Coevolution of human, GenAI, and education
The realm of generative AI evolves as an indispensable tool in the educational sphere. This progression, however, underscores a nuanced balance: generative AI’s development must be aligned with, but not overshadow, the broader trajectory of human advancement. Generative AI should serve as an augmentative force, enhancing human capabilities rather than supplanting them. The essence of this coevolution lies in the acknowledgment that generative AI, while a profound enabler, should not assume a dominant or leadership role. Generative AI’s integration into educational systems should amplify human potential without dictating the course of our development. As shown in Fig. 5, we consider the above issues from the following perspectives.
@@ 5 Discussion
Our response to this inquiry can be succinctly encapsulated from two angles: (1) adaptive learning in the educational sphere has a well-established history spanning several decades. Its efficacy in aiding learners to attain enhanced educational outcomes has been consistently validated through numerous studies. By integrating generative AI with adaptive learning, our goal is to maintain and accelerate the evolution of adaptive learning, propelling it into a new era of advancement; (2) the inherently data-driven nature of adaptive learning fosters the widespread adoption of machine learning algorithms within this field. Drawing on insights and methodologies from other ML-centric areas, such as NLP and CV, we are confident that generative AI will serve as a potent catalyst, steering adaptive learning in novel and exciting directions.
@@
Are all components of adaptive learning compatible with generative AI?
While we’ve highlighted the substantial potential of integrating GenAI into adaptive learning, it’s important to acknowledge the current limitations of GenAI in addressing complex problems. The instructional module of adaptive learning, involving intricate factors like learning theory, pedagogical expertise, and psychological analysis, presents challenges that even domain experts find demanding. Therefore, the next generation of GenAI-driven AL systems will likely still necessitate the nuanced insights of human educators.
@@
What possible essence can GenAI gain from the education realm?
With the advancement of GenAI models demonstrating human-like proficiency in specific tasks, we can harness learning theories originally devised for humans to enhance GenAI’s efficiency. Recent research supports the viability of this approach (Borges et al., 2023). By applying GenAI in developing sophisticated adaptive learning methods, we gain deeper insights into human learning processes. This knowledge can provide innovative ideas for refining GenAI systems, creating a symbiotic relationship between artificial intelligence development and educational theory.
@@ 6 Conclusion
Impact Statements
Our work primarily focuses on advocating for the advancement of educational methods for academic purposes. However, it also brings to light the potential ethical implications of such advancements. Within our discussions, we emphasize crucial issues like fairness, along with the challenges and limitations that accompany these technologies. Although we do not anticipate immediate societal consequences that demand specific attention, we recognize the need for those continuous considerations.

### v3

@@ 4 Impacts of GenAI on AL
In the following, we first outline advantages GenAI that can offer to AL, along with examples of industrial applications (Appendix. A). Then, we provide insights from an educational perspective, delving into the disadvantages and challenges faced by the new framework.
@@ 4.1.1 Diversity and Dynamics
Integrating GenAI into AL offers significant benefits, notably its ability to produce dynamic and diverse outputs. Current ML-based adaptive learning systems, despite their automation (Wang et al., 2023b), still fall short of replicating the nuanced interactions of a skilled human teacher. GenAI bridges this gap by generating real-time, tailored responses and learning plans based on each student’s immediate questions and learning status. Moreover, the recent advancements in LLMs for context understanding (Brown et al., 2020) further enhance the potential for providing personalized feedback, offering a substantial leap in optimizing learning outcomes. Apart from providing new features, GenAI holds the potential to enhance current ML models in AL systems. As detailed in Section 3.1, profile building methods typically use learner’s question practice records, but these are often pre-set and designed for mainstream students, potentially hindering accurate assessment of individual learning status. GenAI, with its dynamic and diverse capabilities, can generate questions tailored to each learner’s past performance, aiding in more effective profile building.
@@ 4.1.2 Multi-modality Convenience
The powerful multi-modal processing capabilities enables GenAI to understand student’s learning status in AL from a broader perspective. To be specific, by integrating GenAI, the AL systems will be able to interpret leaner emotions in speech and facial expressions together with the text input. The enhanced input information will help the following analysis of the student’s learning characteristics such as learning style and learning status (Giannakos & Cukurova, 2023), which will also benefit the precise adaptions to the student’s learning paths. Meanwhile, as GenAI’s outputs can be multi-modal (Wu et al., 2023b), the generated learning materials for knowledge concepts will become more inspiring. For example, with image generation models, the interpretations of geometry problems could be presented with figures or even a short animation, which will be easier for learners to understand compared to prior pure-text responses.
4.1.3 Powerful Prior Knowledge
The scarcity of high-quality annotated data in education is the challenge faced by many ML methods in AL system. And tt is caused by the strong privacy properties and high annotation costs of pedagogical data. The appearance of GenAI models brings new solutions as GenAI models are proven to be general-purpose foundation models (Brühl, 2023). Contributing to the powerful prior knowledge and (Dang et al., 2022) and in-context learning capability (Dong et al., 2022), existing algorithms in AL system are expected to receive a performance boost after the integration it with the prior knowledge provided by GenAI. Besides that, some new tasks with very few annotated data will also be able to get launched through GenAI’s zero-shot or few-shot learning algorithms (Brown et al., 2020). This change will encourages the further explorations of using ML to solve more challenging tasks in AL. At last, the emergent planning prior knowledge enables GenAI not only be a specific executor but a general task planner. It open a door for developing intelligent agent in solving problem through the elaboration of external tools and resources. On the basis of that, some previous complicated tasks in AL will get solved soon.
@@ 4.2.1 Hallucination
The biggest issue with GenAI currently is its potential to produce content that doesn’t actually exist, a phenomenon known as ”Hallucination” (Ji et al., 2023). Hallucinations happened even in the most advanced GenAI models. Research suggests that educating students about the limitations of using GenAI is crucial, emphasizing its susceptibility to errors when working with it (Strzelecki & ElArabawy, 2024). This brings forth numerous opportunities and research questions for the future, such as how to design the AL system to help students cultivate an appropriate reliance on AI-generated content (Vereschak et al., 2021). Furthermore, hallucination could become a new adaptive teaching method, such as creating new and possibly erroneous educational content and asking students to practice the role of a tutor in correcting AI’s mistakes (Ma et al., 2023b).
@@ 4.2.2 Capability Decay
The ease of obtaining information from GenAI can create a dependency, where students might prefer quick GenAI answers over the process of learning and discovery. This over-reliance could erode critical thinking skills and reduce the students’ capacity for self-driven inquiry. One potential solution is to integrate GenAI as a tool that prompts further questioning and exploration rather than providing definitive answers. Encouraging projects and assignments that require independent research and critical analysis can also help maintain a balance between utilizing GenAI and fostering self-reliance. However, these suggestions only begin to address the broader challenge of integrating GenAI into education without compromising the development of independent learning skills. This issue remains an open question, presenting a wide array of opportunities for educators, technologists, and policymakers to explore innovative solutions that balance the benefits of GenAI with the imperative of nurturing autonomous, curious learners.
@@ 4.2.3 Fairness
Fairness is a crucial ethical concern, particularly in education where GenAI technologies are being integrated (Fenu et al., 2022). These advancements, while beneficial, bring to light issues of equitable access and unbiased treatment among learners. For instance, discrepancies between those accessing advanced systems like GPT-4 and those with older versions like GPT-3.5 can lead to unequal knowledge distribution, potentially exacerbating educational gaps. The effective deployment of GenAI in education also hinges on user familiarity and accessibility. Students well-acquainted with AI technologies stand to gain more, while those less experienced may encounter challenges and a steeper learning curve. Introducing standardized training programs might help bridge this gap, though such measures require meticulous implementation and oversight. Additionally, inherent biases within GenAI pose a significant challenge. These systems, limited by the biases in their training data, might offer skewed guidance, making vigilant monitoring essential to ensure fairness. In adaptive learning, where GenAI tailors instructional guidance, the risk of bias amplifies. Personalizing learning experiences is beneficial but needs to be weighed against the risk of unfairness. Ensuring an optimal, equitable, and unbiased educational experience for every learner involves navigating these complexities, making the balance between personalized learning and fairness both a technical hurdle and an ethical necessity in the integration of GenAI into educational systems.
@@ 4.2.4 Coevolution of human, GenAI, and education
The realm of GenAI evolves as an indispensable tool in the educational sphere. This progression, however, underscores a nuanced balance: GenAI’s development must be aligned with, but not overshadow, the broader trajectory of human advancement. Generative AI should serve as an augmentative force, enhancing human capabilities rather than supplanting them. The essence of this coevolution lies in the acknowledgment that GenAI, while a profound enabler, should not assume a dominant or leadership role. Generative AI’s integration into educational systems should amplify human potential without dictating the course of our development. As shown in Fig. 7, we consider the above issues from the following perspectives.
@@ 5 Further Discussion
Our response to this inquiry can be succinctly encapsulated from two angles: (1) adaptive learning in the educational sphere has a well-established history spanning several decades. Its efficacy in aiding learners to attain enhanced educational outcomes has been consistently validated through numerous studies. By integrating GenAI with adaptive learning, our goal is to maintain and accelerate the evolution of adaptive learning, propelling it into a new era of advancement; (2) the inherently data-driven nature of adaptive learning fosters the widespread adoption of machine learning algorithms within this field. Drawing on insights and methodologies from other ML-centric areas, such as NLP and CV, we are confident that GenAI will serve as a potent catalyst, steering adaptive learning in novel and exciting directions.
@@ 6 Conclusion
Authors’ Contributions
Hang L. and QingSong W. conceptualized the manuscript. Hang L. and Tianlong X. created all figures and data visualizations. Hang L., Tianlong X., ChaoLi Z., Eason C. contributed to writing the manuscript, Qingsong W., Jiliang T., Jing L., Xing F., and Haoyang L. contribute to reviewing and editing the manuscript.
Competing interests
The author declares no competing interests.